OpenStudy (anonymous):

If the graph of the quadratic function f(x) = x^2 + dx +3d has its vertex on the x axis what are the possible values of d? What if f(x) = x^2 + 3dx +1?

OpenStudy (anonymous):

complete the square for x^2 + dx +3d to get its vertex

OpenStudy (anonymous):

mukushla is right :) try doing ...

OpenStudy (anonymous):

I have the answers. Easily accessible in the back of the book. How to solve it is the problem. When you say complete the square what do you mean? I've been trying x^2 + dx +3d = ax^2 + (-2ah)x+ (ah^2+k) And trying to identify the vertex.. that waayy... But what do you mean by complete the square?

OpenStudy (anonymous):

the vertex of quadratic equation is given by (-b/2a, c-b^2/4a) since it has vertex on X axis therefore y ordinate =0 . hence c-b^2=0

OpenStudy (anonymous):

Thanks! I'll try it again.

OpenStudy (anonymous):

in 1st case the value of d=0 and 3 in 2nd case d=1/3 and -1/3

OpenStudy (anonymous):

@jenejij take a look http://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (anonymous):

Oh okay. I know how to "complete the square" haha. Got it. =)

OpenStudy (anonymous):

Thanks everyone =)

OpenStudy (anonymous):

Actually - still having problems.. =/ I don't think x=3. x= 0, and x= 12.

OpenStudy (anonymous):

x=12 is not right answer

OpenStudy (anonymous):

thats what the answer in the back of my text book is... but I don't know how to get there.

OpenStudy (anonymous):

what is the final answer of first part?

OpenStudy (anonymous):

d = 0 and d= 12. According the the textbook...

OpenStudy (anonymous):

and for second part ??

OpenStudy (anonymous):

\[d = + or - 2 / \sqrt{13}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I got the answer: plz check it

OpenStudy (anonymous):

the value of x coordinate of vertex is given by -b/2a; from the given question we have a=1 b=d c=3d substituting the values in -b/2a we have -d/2

OpenStudy (anonymous):

to find y ordinate substitute this value in the given eq. we have